# Gauge-independent gravitational waves from a minimal dark $U(1)$ sector with viable dark matter candidates

### v1

@@ 4.1 The tunneling action and transition rate
As the temperature falls below the critical temperature $T_{c}$, the effective potential may develop two local minima separated by a barrier, rendering the minimum in the symmetric phase metastable. The transition from the false vacuum to the true vacuum then proceeds via bubble nucleation. The resulting bubbles expand and collide, sourcing a stochastic gravitational wave background. At finite temperature, the bubble nucleation rate, also referred to as the transition rate or the false vacuum decay rate, is given by [78, 79]
@@
where $S_{3}(T)\equiv S_{3}(\phi_{b},T)$ is the three-dimensional Euclidean action evaluated on the $O(3)$ symmetric bounce configuration. The action is defined as
$S_{3}(\phi,T)=\int_{0}^{\infty}{\rm d}^{3}x\,\bigl[\tfrac{1}{2}(\nabla\phi)^{2}+\widetilde{V}_{{\rm eff}}(\phi,T)\bigr]\,$, (4.2)
with $\widetilde{V}_{{\rm eff}}(\phi,T)\equiv V_{{\rm eff}}(\phi,T)-V_{{\rm eff}}(\phi_{{\rm FV}},T)$ the free energy density shifted relative to the false vacuum. The bounce $\phi_{b}$ satisfies the equation of motion
@@
and subjects to the boundary conditions $\phi(r\to\infty)=0$ and ${\rm d}\phi/{\rm d}r\bigl|_{r=0}=0$, with $r^{2}=\bm{x}^{2}$. In this work the solution is obtained using CosmoTransitions [81].
@@ 4.2 Characteristic temperatures
To ensure the phase transition completes successfully, two conditions must be checked [85, 60]: (1) the transition has an end temperature $T_{\rm e}$, defined by $P_{f}(T_{\rm e})=\varepsilon\lesssim 0.01$; (2) the physical volume of the false vacuum $\mathcal{V}_{{\rm phys}}(t)=a^{3}(t)P_{f}(t)$ decreases with time,
@@
and this condition should be verified at both $T_{p}$ and $T_{\rm e}$.
@@
where $\alpha(T)$ is the transition strength at temperature $T$. For moderate supercooling ($\alpha\lesssim 1$), one has $T_{\rm reh}\approx T_{p}$; whereas for strong supercooling ($\alpha\gg 1$), the reheating temperature can be significantly higher than $T_{p}$, or even higher than $T_{c}$ [60]. The present gravitational wave amplitude and frequency are obtained by redshifting their values from the reheating temperature $T_{\rm reh}$ rather than from $T_{p}$ down to today.
@@
$T_{c}>T_{n}\gg T_{p}\gtrsim T_{\rm e}\gtrsim T_{0}\gtrsim 0\,.$ (4.25)
Recall that $T_{0}$ is defined as the temperature at which $\phi=0$ ceases to be a local minimum. In a typical supercooled phase transition, $T_{p}$, $T_{\rm e}$, and $T_{0}$ are close to one another.
@@ 4.3 Hydrodynamic parameters
The strength of the phase transition is parameterized by $\alpha$, defined as [88, 89]
$\alpha\equiv\frac{\bar{\theta}_{f}(T_{p})-\bar{\theta}_{t}(T_{p})}{3w_{f}(T_{p})}\,,\qquad\bar{\theta}\equiv e-\frac{p}{c_{s,t}^{2}}\,$, (4.30)
where $w_{f}$ denotes the enthalpy density in the false vacuum, $\bar{\theta}$ represents the pseudotrace that is a generalization of the vacuum energy, and $c_{s}^{2}$ is the speed of sound in the plasma. Subscripts $f$ and $t$ refer to quantities in the false and true vacua respectively. The hydrodynamics of the plasma is described by the energy density $e$, pressure $p$, and enthalpy density $w$. The plasma is usually modeled as a perfect relativistic fluid, for which the pressure is related to the free energy density $\mathcal{F}$ by $p=$-$\mathcal{F}$. Since $\mathcal{F}$ equals to the effective potential, one has $p=-V_{{\rm eff}}$. The relevant thermodynamic quantities are given by
@@
$e$ $\equiv T\frac{\partial p}{\partial T}-p\,$, (4.32)
$w$ $\equiv T\frac{\partial p}{\partial T}=p+e\,$, (4.33)
$c_{s}^{2}$ $\equiv\frac{\delta p}{\delta e}\simeq\frac{\partial p/\partial T}{\partial e/\partial T}\,$, (4.34)
where the second term in $p$ accounts for the field-independent contribution from all relativistic particle species in thermal equilibrium with the dark Higgs $h_{x}$ at temperature $T$ [89].
@@
Kinetic energy fraction is defined as $K\equiv\rho_{{\rm fl}}/e_{p}$ which quantifies the fraction of available vacuum energy converted into bulk motion of the plasma, where $e_{p}$ is the total energy density at the percolation temperature $T_{p}$, and $\rho_{{\rm fl}}$ is the plasma kinetic energy density, obtained by integrating over the fluid profile. The kinetic energy fraction can be further parameterized as
@@ 4.4 Gravitational wave power spectrum
where $f$ denotes the frequency and $\rho_{c}$ is the critical density. The total gravitational wave signal typically receives three primary contributions: (1) bubble collisions of the scalar field; (2) sound waves in the bulk plasma; (3) magneto-hydrodynamic turbulence in the plasma [93, 94, 95]. Accordingly,
@@
Finally, noted from gravitational wave power spectrum formulas, the peak frequency parametrically scales as
@@ 4.5.1 Gauge-dependent gravitational wave results
Although results obtained from a gauge-dependent potential do not constitute definitive gravitational wave predictions for the $U(1)$ dark sector, the two panels nevertheless reveal clear trends that are instructive for understanding the underlying physics: larger symmetry breaking scales $v_{x}$ shift the gravitational wave signal toward higher frequencies, while smaller $v_{x}$ shift it toward lower frequencies. In our scan, the lowest peak frequencies approach the nanohertz band with $v_{x}\lesssim 100$ MeV.
The temperature at which the transition occurs is also closely correlated with the signal strength: low-temperature transitions typically yield larger amplitudes, whereas transitions occurring at higher temperatures tend to produce much weaker signals.
@@
We identify an approximately high-temperature regime using the criterion $T_{p}/m_{A^{\prime}}\gtrsim 2$, which already constitutes a small subset of the successful points (below $5\%$). Truly high-temperature transitions for which the high-temperature expansion is parametrically justified, $T_{p}/m_{A^{\prime}}\gg 10$, are rare in our scan. These points generally produce very weak signals and are therefore difficult to detect, making them phenomenologically less compelling.
@@
It should be noted that part of the literature adopts Eq. (3.57) as the starting point for gravitational wave analyses and reports many viable points in the intermediate- and low-temperature regimes. Such treatments are not appropriate. Eq. (3.57) is obtained by employing the high-temperature expansion of the $J_{B}$ functions, which ceases to be valid once the high-temperature condition is no longer satisfied. Consequently, the resulting gravitational wave spectra in regimes where the high-temperature expansion is not valid are unreliable, regardless of gauge-dependent issues.
@@ 5.1.2 Dark fermion dark matter from the $U(1)$ portal
The benchmark models we consider are chosen such that the dark Higgs $h_{x}$ remains in thermal equilibrium and decays into SM fermions before BBN, and thus without affecting the number density evolution of the dark matter or the dark photon. We therefore focus on the coupled evolution of $\chi$ and $A^{\prime}$. The dominant annihilation channel for dark fermion dark matter is
@@ 5.2 Benchmark models
Although these benchmarks satisfy the dark matter relic density constraint, they typically require a tiny dark gauge coupling $g_{x}$ to avoid excessive secluded depletion of the the dark photon. As a result, the phase transition is relatively weak and the corresponding gravitational wave signal is suppressed. Moreover, for these benchmarks we find that $T_{p}/m_{A^{\prime}}$ lies outside the validity ranges of our gauge-independent treatments in both the high- and low-temperature limits.
@@
In model (d), the dark fermion is heavier than the dark photon, $m_{\chi}>m_{A^{\prime}}$. As the Universe cools, $\chi\bar{\chi}$ annihilates efficiently into dark photons and yields the observed relic abundance. The dark photons subsequently decay into $e^{+}e^{-}$ pairs before BBN. For models (e) and (f), we instead choose $m_{\chi}$ to lie slightly below $m_{A^{\prime}}$. This near-degeneracy allows for a larger $g_{x}$ while still reproducing the correct $\Omega_{\chi}h^{2}$ via the process $A^{\prime}A^{\prime}\to\chi\overline{\chi}$. We indicate these three Case 2 benchmarks in Figs. 5 and 7. Benchmark models (e) and (f) lie within the projected sensitivity of Taiji, TianQin, and LISA, and thus provide promising multi-messenger targets for a genuinely minimal $U(1)$ dark sector.
@@
Figure 7: Experimental constraints on the kinetic mixing parameter $\delta$, defined in Eqs. (2.4) and (5.10), for dark photon masses $m_{A^{\prime}}\in[1{\rm MeV},\,100{\rm GeV}]$. The three Case 2 benchmark models (Table 3) are shown as black crosses and are consistent with all current experimental bounds. Constraints on the kinetic mixing for dark photons with $m_{A^{\prime}}\gtrsim$ 1 MeV arise from supernova (SN1987A) [118], $(g-2)_{e}$ [119] and from various beam dump experiments: E141 [120], E137 [121, 122, 123], $\nu$-Cal [124, 125]. The pink region in the upper portion of the plot, labeled as “Others”, collects constraints from collider and fixed-target experiments, including: CMS [126], BaBar [127], LHCb [128], KLOE [129, 130, 131, 132] and A1 [133].
@@ 5.3 Phenomenological implications
At the level of thermal freeze-out, the tension is particularly transparent: with present experimental sensitivities, many WIMP-like candidates that remain compatible with direct detection bounds cannot maintain sufficient annihilation into SM final states, which leads to an overproduction of dark matter. In hidden sector constructions, this difficulty can be even more pronounced, since the interactions connecting the dark sector to the SM are typically mediated by portal couplings that must be small. Consequently, annihilation directly into SM particles can be inefficient. An common alternative is secluded annihilation, in which dark matter annihilates dominantly into lighter dark mediators that subsequently decay into SM states, thereby achieving an efficient depletion of the relic abundance while alleviating dark matter experimental constraints.
@@ 6 Conclusion and outlook
Using the gauge-independent effective action, we scanned the high- and low-temperature limits separately and then retained only those points consistent with the corresponding temperature-limit assumptions. The surviving points shown in Fig. 5 provide concrete, gauge-independent predictions for the minimal $U(1)_{x}$ setup, mapping viable microscopic parameters to detector-facing peak frequencies and amplitudes. A notable outcome is that supercooled transitions constitute a substantial fraction of successful transitions in broad scans and, when realized, often populate the most promising regions of experimental sensitivity, reinforcing the view that supercooled first-order phase transitions are among the most compelling targets in current gravitational wave searches.
By contrast, parametrically high-temperature transitions are comparatively rare and typically yield weak signals. Between these limits lies an intermediate regime in which neither the high- nor the low-temperature gauge-independent treatment is applicable. Bridging this gap with a fully gauge-independent framework remains an important open direction.
@@
Across all temperature regimes, we find a robust feature of the parameter region that supports a first-order phase transition: the dark photon must be heavier than the dark Higgs. If nature were to realize the opposite hierarchy, a $U(1)_{x}$ breaking phase transition in the minimal dark sector and thus the associated gravitational wave signal, will not occur.

### v2

@@ 4.1 The tunneling action and transition rate
As the temperature falls below the critical temperature $T_{c}$, the effective potential may develop two local minima separated by a barrier, rendering the minimum in the symmetric phase metastable. The transition from the false vacuum to the true vacuum then proceeds via bubble nucleation. The resulting bubbles expand and collide, sourcing a stochastic gravitational wave background. At finite temperature, the bubble nucleation rate, also referred to as the transition rate or the false vacuum decay rate, is given by [82, 83, 84]
@@
where $S_{3}(T)\equiv S_{3}(\phi_{b},T)$ is the three-dimensional Euclidean action evaluated on the $O(3)$ symmetric bounce configuration. The action is defined as [^10] At finite temperature $T$, the Euclidean time direction is compactified with period $\beta\equiv 1/T$. The Euclidean action thus takes the form $S_{E}(\phi)=\int_{0}^{\beta}{\rm d}\tau\int{\rm d}^{3}x\left[\tfrac{1}{2}(\partial_{\tau}\phi)^{2}+\tfrac{1}{2}(\nabla\phi)^{2}+V_{\rm eff}(\phi,T,\xi)\right]\,.$ (4.2) In the conventional gauge-dependent treatment one often works in Landau gauge, $\xi=0$. By contrast, in the gauge-independent approach discussed in Section 3.2, we take $V_{\rm eff}$ at leading order, for which it is independent of the gauge-fixing parameter $\xi$, and thus the resulting bounce solution is manifestly gauge independent. For thermally assisted bubble nucleation, the dominant saddle point configuration is approximately time-independent ($\partial_{\tau}\phi\simeq 0$). In this case the action reduces to $S_{E}\simeq\beta\,S_{3}(T)=S_{3}(T)/T\,$, (4.3) where $S_{3}(T)$ is shown in Eq. (4.4).
$S_{3}(\phi,T)=\int_{0}^{\infty}{\rm d}^{3}x\,\left[\frac{1}{2}(\nabla\phi)^{2}+\widetilde{V}_{{\rm eff}}(\phi,T)\right]\,$, (4.4)
with $\widetilde{V}_{{\rm eff}}(\phi,T)\equiv V_{{\rm eff}}(\phi,T)-V_{{\rm eff}}(\phi_{{\rm FV}},T)$ the free energy density shifted relative to the false vacuum. Varying $S_{3}$ yields the three-dimensional bounce equation. With the $O(3)$ symmetry, the bounce depends only on the radial coordinate, $\phi=\phi(r)$ with $r=\sqrt{x^{2}+y^{2}+z^{2}}$,
@@
supplemented by the boundary conditions
$\phi\,(r\to\infty)=0\,,\qquad\left.\frac{{\rm d}\phi}{{\rm d}r}\right|_{r=0}=0\,.$ (4.6)
In this work the solution is obtained using CosmoTransitions [85].
@@ 4.2 Characteristic temperatures
To ensure the phase transition completes successfully, two conditions must be checked [89, 64]: (1) the transition has an end temperature $T_{e}$, defined by $P_{f}(T_{e})=\varepsilon\lesssim 0.01$; (2) the physical volume of the false vacuum $\mathcal{V}_{{\rm phys}}(t)=a^{3}(t)P_{f}(t)$ decreases with time,
@@
and this condition should be verified at both $T_{p}$ and $T_{e}$.
@@
where $\alpha(T)$ is the transition strength at temperature $T$. For moderate supercooling ($\alpha\lesssim 1$), one has $T_{\rm reh}\approx T_{p}$; whereas for strong supercooling ($\alpha\gg 1$), the reheating temperature can be significantly higher than $T_{p}$, or even higher than $T_{c}$ [64]. Our scan indicates that the minimal gauged $U(1)$ dark sector typically does not realize a very strong supercooled phase transition, the corresponding reheating temperature generally does not have $T_{\rm reh}>T_{c}$. The present gravitational wave amplitude and frequency are obtained by redshifting their values from the reheating temperature $T_{\rm reh}$ rather than from $T_{p}$ down to today.
@@
$T_{c}\gg T_{n}>T_{p}\gtrsim T_{e}\gtrsim T_{0}\gtrsim 0\,.$ (4.28)
Recall that $T_{0}$ is defined as the temperature at which $\phi=0$ ceases to be a local minimum. In a typical supercooled phase transition, $T_{p}$, $T_{e}$, and $T_{0}$ are close to one another.
@@ 4.3 Hydrodynamic parameters
The phase transition strength parameter $\alpha$, which quantifies the energy released during the transition relative to the radiation energy density of the Universe at the time the gravitational waves are generated, is defined as [92, 93]
$\alpha\equiv\frac{\bar{\theta}_{f}(T_{p})-\bar{\theta}_{t}(T_{p})}{3w_{f}(T_{p})}\,,\qquad\bar{\theta}\equiv\rho-\frac{p}{c_{s,t}^{2}}\,$, (4.33)
where $w_{f}$ denotes the enthalpy density in the false vacuum, $\bar{\theta}$ represents the pseudotrace that is a generalization of the vacuum energy, and $c_{s}^{2}$ is the speed of sound in the plasma. Subscripts $f$ and $t$ refer to quantities in the false and true vacua respectively. The hydrodynamics of the plasma is described by the energy density $\rho$, the pressure $p$, and the enthalpy density $w$.
The plasma is usually modeled as a perfect relativistic fluid, with the pressure in the false vacuum determined by the negative finite-temperature free energy density. For a dark sector first-order phase transition, the pressure receives two contributions:
  • The dark Higgs sector contribution is encoded in the effective potential $V_{\rm eff}(\phi,T)$, which includes all particles that couple to the dark Higgs and therefore contribute to the $\phi$-dependent thermodynamics.
  • The contribution from additional radiation species that do not couple directly to the dark Higgs and therefore do not contribute directly to the $\phi$-dependent part of $V_{\rm eff}$, but still belong to the dark thermal bath. These species should be included only if they remain relativistic and in kinetic equilibrium with the dark sector at the time of the phase transition. In a realistic $U(1)_{x}$ dark sector, this typically refers to particles that interact efficiently with the dark photon (but not with the dark Higgs), provided that the dark photon itself remains in equilibrium with the dark Higgs.
The relevant thermodynamic quantities are thus given by [93, 38]
@@
$\rho$ $\equiv T\frac{\partial p}{\partial T}-p\,$, (4.35)
$w$ $\equiv T\frac{\partial p}{\partial T}=p+\rho\,$, (4.36)
$c_{s}^{2}$ $\equiv\frac{\delta p}{\delta\rho}\simeq\frac{\partial p/\partial T}{\partial\rho/\partial T}\,$, (4.37)
where the second term in $p$ accounts for the field-independent contribution from all relativistic particle species in thermal equilibrium with the dark thermal bath at temperature TT. [^11] For a $U(1)_{B-L}$ phase transition, all SM fermions carry BLB-L charge. Although the SM fermions are vectorlike under BLB-L and therefore do not couple directly to the $U(1)_{B-L}$ Higgs, they interact efficiently with the $U(1)_{B-L}$ gauge boson and thus remain in thermal equilibrium with the $U(1)_{B-L}$ plasma. Consequently, SM fermions contribute to the second term in Eq. (4.34) provided that they are still relativistic at the time of the $U(1)_{B-L}$ phase transition. This makes the $U(1)_{B-L}$ case qualitatively different from a dark $U(1)_{x}$ sector phase transition, in which SM fermions carry no $U(1)_{x}$ charge.
@@
Kinetic energy fraction is defined as $K\equiv\rho_{{\rm fl}}/\rho_{p}$ which quantifies the fraction of available vacuum energy converted into bulk motion of the plasma, where $\rho_{p}$ is the total energy density at the percolation temperature $T_{p}$, and $\rho_{{\rm fl}}$ is the plasma kinetic energy density, obtained by integrating over the fluid profile. The kinetic energy fraction can be further parameterized as
@@ 4.4 Gravitational wave power spectrum
where $f$ denotes the frequency and $\rho_{c}$ is the critical density.
The total gravitational wave signal typically receives three primary contributions: (1) bubble collisions of the scalar field; (2) sound waves in the bulk plasma; (3) magneto-hydrodynamic turbulence in the plasma [97, 98, 99]. Accordingly,
@@
Finally, as can be seen from gravitational wave power spectrum formulas, the peak frequency parametrically scales as
@@ 4.5.1 Gauge-dependent gravitational wave results
Although results obtained from a gauge-dependent potential do not constitute definitive gravitational wave predictions for the dark $U(1)$ sector, the two panels nevertheless reveal clear trends that are instructive for understanding the underlying physics:
  • The temperature at which the transition occurs is closely correlated with the signal strength: low-temperature transitions typically yield larger amplitudes, whereas transitions occurring at higher temperatures tend to produce much weaker signals.
  • Larger symmetry breaking scales $v_{x}$ shift the gravitational wave signal toward higher frequencies, while smaller $v_{x}$ shift it toward lower frequencies.
@@
We identify an approximately high-temperature regime using the criterion $T_{p}/m_{A^{\prime}}\gtrsim 2$, which already constitutes a small subset of the successful points (below $5\%$). Truly high-temperature transitions for which the high-temperature expansion is parametrically justified, $T_{p}/m_{A^{\prime}}\gtrsim 10$, are rare in our scan. These points generally produce very weak signals and are therefore difficult to detect, making them phenomenologically less compelling.
@@
It should be noted that part of the literature adopts Eq. (3.16) as the starting point for gravitational wave analyses and reports many viable points in the intermediate- and low-temperature regimes. Such treatments are not appropriate. Eq. (3.16) is obtained by employing the high-temperature expansion of $J_{B},J_{F}$ functions, which ceases to be valid once the high-temperature condition is no longer satisfied. Consequently, the resulting gravitational wave spectra in regimes where the high-temperature expansion is not valid are unreliable, regardless of gauge-dependent issues.
@@ 4.5.2 Gauge-independent gravitational wave predictions
As can be seen from Fig. 5, in the minimal $U(1)$ dark sector, stochastic gravitational wave signals within the reach of current and planned detectors arise predominantly from supercooled phase transitions. In particular, $U(1)_{x}$ vev in the range $v_{x}\in[10,100]\penalty 10000\ \mathrm{MeV}$ can yield peak frequencies approaching the nanohertz band, offering a plausible interpretation of the PTA signal region. While for $v_{x}\in[1,100]\penalty 10000\ \mathrm{GeV}$ the predicted peak frequencies fall into the millihertz range targeted by proposed space-based interferometers such as Taiji, TianQin, and LISA, providing a promising discovery window.
@@ 5.1.2 Dark fermion dark matter from the $U(1)$ portal
The benchmark models we consider are chosen such that the dark Higgs $h_{x}$ remains in thermal equilibrium and decays into SM fermions before BBN, thus not affecting the number density evolution of the dark matter or the dark photon. We therefore focus on the coupled evolution of $\chi$ and $A^{\prime}$. The dominant annihilation channel for dark fermion dark matter is
@@ 5.2 Benchmark models
Although these benchmarks satisfy the dark matter relic density constraint, they typically require a tiny dark gauge coupling $g_{x}$ to avoid excessive secluded depletion of the dark photon. As a result, the phase transition is relatively weak and the corresponding gravitational wave signal is suppressed. Moreover, for these benchmarks we find that $T_{p}/m_{A^{\prime}}$ lies outside the validity ranges of our gauge-independent treatments in both the high- and low-temperature limits.
@@
In model (d), the dark fermion is heavier than the dark photon, $m_{\chi}>m_{A^{\prime}}$. As the Universe cools, $\chi\bar{\chi}$ annihilates efficiently into dark photons and yields the observed relic abundance. The dark photons subsequently decay into $e^{+}e^{-}$ pairs before BBN. For models (e) and (f), we instead choose $m_{\chi}$ to lie slightly below $m_{A^{\prime}}$. This near-degeneracy allows for a larger $g_{x}$ while still reproducing the correct $\Omega_{\chi}h^{2}$ via the process $A^{\prime}A^{\prime}\to\chi\overline{\chi}$. We indicate these three Case 2 benchmarks in Figs. 5 and 7. Benchmark models (e) and (f) lie within the projected sensitivity of Taiji, TianQin, and LISA, and thus provide promising multi-messenger targets for a genuinely minimal gauged $U(1)$ dark sector.
@@
Figure 7: Experimental constraints on the kinetic mixing parameter $\delta$, defined in Eqs. (2.4) and (5.10), for dark photon masses $m_{A^{\prime}}\in[1{\rm MeV},\,100{\rm GeV}]$. The three Case 2 benchmark models (Table 3) are shown as black crosses and are consistent with all current experimental bounds. Constraints on the kinetic mixing for dark photons with $m_{A^{\prime}}\gtrsim$ 1 MeV arise from supernova (SN1987A) [122, 123], $(g-2)_{e}$ [124] and from various beam dump experiments: E141 [125], E137 [126, 127, 128], $\nu$-Cal [129, 130]. The pink region in the upper portion of the plot, labeled as “Others”, collects constraints from collider and fixed-target experiments, including: CMS [131], BaBar [132], LHCb [133], KLOE [134, 135, 136, 137] and A1 [138].
@@ 5.3 Phenomenological implications
At the level of thermal freeze-out, the tension is particularly transparent: with present experimental sensitivities, many WIMP-like candidates that remain compatible with direct detection bounds cannot maintain sufficient annihilation into SM final states, which leads to an overproduction of dark matter. In hidden sector constructions, this difficulty can be even more pronounced, since the interactions connecting the dark sector to the SM are typically mediated by portal couplings that must be small. Consequently, annihilation directly into SM particles can be inefficient. A common alternative is secluded annihilation, in which dark matter annihilates dominantly into lighter dark mediators that subsequently decay into SM states, thereby achieving an efficient depletion of the relic abundance while alleviating dark matter experimental constraints.
@@ 6 Conclusion and outlook
Using the gauge-independent effective action, we scanned the high- and low-temperature limits separately and retained only those parameter points that are self-consistent with the corresponding temperature-limit assumptions. The surviving points shown in Fig. 5 hence provide concrete, gauge-independent predictions for the minimal gauged $U(1)_{x}$ dark sector, directly mapping viable microscopic parameters to detector-facing peak frequencies and amplitudes. We find that, in the minimal $U(1)_{x}$ dark sector, detectable signals in current and planned experiments arise predominantly from supercooled phase transitions. In particular, a $U(1)_{x}$ symmetry breaking scale $v_{x}\in[10,100]\penalty 10000\ \mathrm{MeV}$ can yield peak frequencies approaching the nanohertz band, overlapping the PTA signal region; whereas $v_{x}\in[1,100]\penalty 10000\ \mathrm{GeV}$ leads to peak frequencies in the millihertz range targeted by proposed space-based interferometers such as Taiji, TianQin, and LISA.
A notable outcome of our scan is that supercooled transitions make up a sizable fraction of viable points and, when realized, tend to populate the most sensitive regions of the experimental landscape. By contrast, parametrically high-temperature transitions are comparatively rare and typically produce weaker signals. Between these limits lies an intermediate regime in which neither the high- nor low-temperature gauge-independent treatment is applicable. Bridging this gap with a fully gauge-independent framework remains an important open direction.
@@
Across all temperature regimes, we find a robust feature of the parameter region that supports a first-order phase transition: the dark photon must be heavier than the dark Higgs. If nature were to realize the opposite hierarchy, a $U(1)_{x}$ breaking phase transition in the minimal dark sector and thus the associated gravitational wave signal will not occur.